OpenStudy (anonymous):

HELP!! The center of a circle is at (5, 12) and it has a radius of 16. What is the equation of the circle? (x + 5)2 + (y 12)2 = 256 (x + 5)2 + (y 12)2 = 4 (x 5)2 + (y + 12)2 = 256 (x 5)2 + (y + 12)2 = 4

OpenStudy (zale101):

you know what's the equation for the circle?

OpenStudy (anonymous):

No i dont know anything pls help me with this question

OpenStudy (anonymous):

the equation for a circle is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

so would it be the third one?

OpenStudy (zale101):

(x-a)2 + (y-b)2 = r^2 Josh already motioned it,

OpenStudy (anonymous):

I think they are all wrong because if you plug in those numbers it will actually be -a and -b

OpenStudy (anonymous):

its one of them because it multiple chocie

OpenStudy (anonymous):

mm are you sure that in the question theres a negative?

OpenStudy (anonymous):

The center of a circle is at (-5, 12) and it has a radius of 16. What is the equation of the circle?

OpenStudy (zale101):